# industrial control system STIX Structured Threat Information Expression

### v1

@@ 1 Introduction
As critical infrastructure increasingly becomes the target of cyber threats, organizations and governments must be able to detect and share information about threat actor activity. Numerous government reports, such as Executive Order 13691 [exec_order_13691] and the Cyberspace Solarium Commission report [cyberspace_solarium_2023], have highlighted the need for enhanced information sharing to safeguard critical infrastructure. Additionally, numerous government programs have been established to facilitate information sharing between the public and private sectors.
@@
In this paper, we examine the challenges that hinder the sharing of actionable threat information with asset owners within the ICS domain. This study leverages the MITRE ATT&CK for ICS [mitre_attack_ics] knowledge base, which aggregates cyber threat intelligence reports across known ICS incidents.
Our paper highlights the key challenges by reviewing three well-known incidents: Stuxnet [stuxnet_facts_report], Industroyer [ukraine_power_grid_analysis], and Triton [saudi_arabia_cyberattack]. This analysis aims to identify the adversarial techniques employed by threat actors and to determine the information that asset owners need to implement effective detection mechanisms against these techniques. We also explore nine recent ICS vulnerability advisories with evidence of exploitation as documented in the CISA Known Known Exploitable Vulnerability (KEV) Catalog  [cisa_kev_catalog].
@@ 2 Related Work
Recent advancements in Cyber Threat Intelligence (CTI) have underscored the complexities of securing ICS and broader Cyber-Physical Systems, particularly in the context of critical infrastructure. ICSs and CPSs often rely on proprietary and specialized protocols that are difficult to monitor and integrate with conventional cybersecurity tools. Krasznay et al. [krasznay2021possibilities] highlighted the critical vulnerabilities in these systems, which stem from their original design priorities focused on safety rather than cybersecurity. The increasing connectivity of ICS and Operational Technology (OT) systems has exposed these vulnerabilities to sophisticated adversaries, including state-sponsored actors, necessitating a paradigm shift in cybersecurity practices. The SeConSys initiative in Hungary represents a significant milestone in addressing these challenges. This initiative facilitates sector-specific threat intelligence sharing and detection through the development of tailored CTI feeds, honeypots that simulate energy-sector environments, and the implementation of the Network and Information Systems (NIS) Directive. These efforts have been instrumental in promoting localized, collaborative approaches to cybersecurity and advancing frameworks for real-time threat information sharing among key stakeholders in the energy sector.
López-Morales et al. [lopez2024securing] extended these contributions by developing innovative methodologies that advance the state-of-the-art in CTI for CPSs. Their work introduces the ICS2 Matrix, an enhanced version of the MITRE ATT&CK framework tailored specifically to ICSs. The ICS2 Matrix incorporates novel attack techniques and mitigation strategies derived from an extensive analysis of Programmable Logic Controller (PLC)-related security research spanning 17 years. Additionally, the development of a high-interaction satellite honeypot provides a groundbreaking tool for simulating realistic adversary interactions, collecting data on emerging threats, and advancing the understanding of satellite-specific attack vectors. This work further proposes the creation of a sandbox environment for connected autonomous vehicles, enabling the simulation and analysis of cyberattacks on autonomous systems to improve resilience. Such contributions underscore the necessity of domain-specific CTI methodologies that address the unique operational and security challenges inherent in different CPSs.
@@ 2.1 Empirical Measurements and Quality Assessment of CTI Sharing
At the ecosystem level, empirical studies have begun to quantify the effectiveness and challenges of CTI sharing at scale. Jin et al. [jin2024sharing] provided one of the first comprehensive empirical measurements of CTI sharing effectiveness, analyzing STIX volume, timeliness, coverage, and quality across public sources. Their findings reveal significant variations in feed quality and highlight the challenges that organizations encounter while operating in a shared intelligence environment. This empirical foundation complements broader systematic reviews that examine how organizations actually use CTI for security decision-making [ainslie2023cyber] and identify persistent barriers to effective CTI sharing, including governance issues, misaligned incentives, and insufficient standardization [alaeifar2024current].
Quality and trust of shared feeds remain persistent issues across the CTI ecosystem. Griffioen et al. [griffioen2019quality] conducted systematic quality evaluations of cyber threat intelligence feeds, revealing inconsistent timeliness and specificity that can undermine the operational value of shared intelligence. Similarly, Schaberreiter et al. [schaberreiter2019quantitative] developed quantitative frameworks for evaluating trust in CTI quality, demonstrating that trust and quality metrics vary significantly across different sources and feed types. Together, these studies highlight the need for more rigorous quality assurance mechanisms in CTI sharing initiatives.
@@ 2.2 Standardization and Technical Implementation
Standardization efforts play a crucial role in making threat observables portable and actionable across different organizational contexts. STIX 2.1, approved as an OASIS Standard (on 10 June 2021), serves as the exchange format used across many CTI producers and consumers [stix_v21_official]. However, challenges remain in effectively representing complex ICS-specific artifacts within existing STIX objects. Zych and Mavroeidis [zych2022enhancing] have proposed enhancements to MITRE ATT&CK representations by STIX, to improve filtering and prioritization tasks, while RFC 9424 [rfc9424] provides important clarification on the capabilities and limitations of IOCs, reinforcing the value of higher-fidelity observables for creating durable detection capabilities.
@@
Recent advances in natural language processing have introduced new approaches for extracting threat intelligence from unstructured reports. While some pipelines pursue full STIX coverage through classical NLP techniques combined with external knowledge bases [marchiori2023stixnet], alternative approaches rely on large language models to extract observables directly from text without requiring extensive knowledge graphs. These automated extraction techniques show promise for scaling CTI processing, though systematic evaluation reveals significant limitations and failure modes. Alam et al. [alam2024ctibench] developed CTIBench, a comprehensive benchmark specifically designed to evaluate LLM performance in cyber threat intelligence tasks across multiple cognitive dimensions, including memorization, understanding, problem-solving, and reasoning. Their evaluation of five state-of-the-art models revealed concerning patterns: all models struggled significantly with tasks requiring domain expertise, with even the best-performing model (GPT-4) achieving only 71% accuracy on foundational CTI knowledge questions. More critically, their analysis identified systematic failure modes, including overestimation of threat severity scores, difficulty with out-of-knowledge cutoff information, and particularly poor performance on questions related to mitigation strategies and adversarial tools. These findings highlight substantial gaps in LLM reliability for cybersecurity applications.
@@
Fieblinger et al. [fieblinger2024actionable] developed a comprehensive methodology for automating the extraction of actionable CTI using LLMs for generating Knowledge Graphs. Their approach systematically evaluates multiple open-source LLMs (including Llama 2 series, Mistral 7B Instruct, and Zephyr) across different extraction techniques, such as few-shot prompt engineering, guidance frameworks, and fine-tuning to optimize the extraction of meaningful triples from CTI texts. Notably, their methodology demonstrates that guidance frameworks significantly improve extraction performance beyond what is achievable with prompt engineering alone, while fine-tuning proves essential for extracting triples that adhere to specified ontologies. The extracted triples are subsequently used to construct knowledge graphs that provide structured, queryable representations of threat intelligence, facilitating downstream applications such as link prediction and threat analysis.
@@
The development of domain-specific evaluation datasets has become increasingly important for assessing automated threat intelligence processing. Bhusal et al. [bhusal2024secure] developed the SECURE benchmark using MITRE ATT&CK as a foundation for generating cybersecurity LLM performance evaluation across extraction, understanding, and reasoning tasks. Their approach demonstrates the potential of leveraging structured threat intelligence frameworks like ATT&CK for creating systematic evaluation datasets. However, the reliability of LLM-based approaches in cybersecurity contexts remains a critical concern. Their comprehensive evaluation of seven state-of-the-art models revealed significant limitations, including hallucinations when models lack current information, difficulty with out-of-distribution detection, and substantial performance variations across different cybersecurity task types. Importantly, their findings demonstrate that while LLMs show promise for cybersecurity advisory roles, closed-source models consistently outperform open-source alternatives, particularly in problem-solving and out-of-distribution scenarios. These results underscore the importance of meticulous validation and human oversight when deploying LLMs for threat intelligence processing, particularly in critical infrastructure environments where inaccurate information can have severe consequences.
2.2.1 ATT&CK Framework Application in OT Environments
@@
For technique-level grounding in OT environments, the MITRE ATT&CK framework and its ICS extension provide structured approaches to understanding adversary behaviors [mitre_attack_design, mitre_attack_ics_design]. Afenu et al. [afenu2024industrial] have leveraged ATT&CK to validate ICS defense mechanisms and testbed implementations. Choi et al. [choi2021probabilistic] generated realistic OT attack sequences based on analysis of real-world datasets. These applications demonstrate the framework’s utility in bridging the gap between theoretical threat models and practical security implementations.
@@
The existing literature highlights several recurring challenges in CTI, including limited interoperability of detection tools, insufficient standardization of threat-sharing protocols, and a lack of actionable intelligence for specific adversarial behaviors. While Krasznay et al. [krasznay2021possibilities] emphasized the importance of legislative and collaborative frameworks, López-Morales et al. [lopez2024securing] focused on advancing technical methodologies to enhance the granularity and applicability of CTI. Both approaches converge on the need for tailored threat intelligence that integrates sector-specific requirements into practical and effective security solutions. Building on these foundational works, this study aims to address critical gaps in the ICS threat information-sharing ecosystem. Our research specifically focuses on expanding the cyber observable objects schema within the STIX standard to ensure the accurate representation of ICS-specific artifacts and adversarial techniques. Unlike prior studies that primarily concentrated on domain-specific or geographically localized advancements, this work seeks to generalize CTI enhancements across diverse ICS environments. By integrating detailed technical insights into standardized threat-sharing protocols, this study aspires to bridge the gap between theoretical threat models and actionable intelligence, thereby facilitating more effective collaboration and situational awareness across ICS stakeholders.
@@ 3 Key Roles and Functions in Sharing Threat/Vulnerability Information
Typically, governments facilitate information sharing by providing reports with threat or vulnerability information to critical infrastructures and other organizations [gao_cybersecurity_2023]. The National Cyber Emergency Response Teams (CERT) is an important information-sharing hub, as it has unique access to threat information through the facilitation of information-sharing programs with both federal, defense, and intelligence agencies, as well as partnerships with private sector organizations.
Security vendors provide tools/platforms and information to support the detection and mitigation of cyber threats [splunk_tdir, crowdstrike_tdir]. Many of these vendors search for vulnerabilities and threat actor activity to develop analytics or rules to be incorporated into detection tools. Some vendors provide machine-readable threat information feeds, which can be directly consumed by security products, including those used to monitor for cyber threats. Further, many vendors obtain data from their detection infrastructures deployed at asset owner sites or collect information while supporting incident response efforts.
ICS vendors typically support information sharing efforts through the disclosure of security advisories and patches when a vulnerability is discovered in their software or devices [ics_patch_tuesday]. Most importantly, ICS vendors have unique knowledge regarding the technical details for their products, which are typically based on proprietary technology.
Asset owners are responsible for managing the risk presented by the cyber threats. The larger organizations that maintain critical infrastructures typically have dedicated security operations teams, the smaller commercial asset owners commonly rely on their national CERT [Cybersecurity_Guide_for_SMEs]. They may have a person partially employed in security operations to comply with regulations. The security team of an asset owner is responsible for monitoring their systems for threats and reporting to their auditors or the CERT.
@@
  • Standardized threat information languages: One key contribution to make threat information sharing more effective is the development of standard languages. For example, STIX standard provides a JSON format to communicate threat behaviors. This includes definitions of SCOs [stix_v21_official] for specific types of artifacts that can be used to detect malicious activities or collect information about them. Other formats include OpenIOC [^1] https://www.scribd.com/document/653654203/An-Introduction-to-OpenIOC and MISP XML [^2] https://www.misp-project.org/.
@@ 4.1.1 Triton
The Triton malware [CISA2017HatMan], also known as TRISIS or HatMan, targeted safety instrumented systems within an oil refinery by manipulating both an engineering workstation and the Triconex safety controllers used in the sulfur recovery units and burn management systems. While the broader campaign during this attack included activities used to gain access to the facility (references as TEMP.Veles in ATT&CK), the defined Tactics, Techniques and Proceduress specific to the ICS environment primarily address the behaviors associated with the manipulation of Windows-based engineering workstations, safety controllers, and the associated network communications between them. The following sections will review the adversary activity associated with these three areas based on the previously defined criteria.
@@
TriStation Protocol (Table 2): The TriStation protocol, which is a proprietary protocol without any available open-source parsers, was heavily used throughout the attack. The adversarial techniques using this protocol include discovering target devices [T0846], detecting and changing their operating mode [T0858] [T0868], downloading programs to the targeted safety controller [T0843], executing the program [T0871], and extracting programs from the device [T0845]. Developing detections for this activity requires the ability to (i) understand and parse the TriStation protocol  [tristation_protocol_reversing], (ii) identify specific functions of the protocol, and (iii) represent function payloads (e.g., program downloads contents). The STIX Network Traffic observable could be used to represent this activity, specifically the src/dst_payload_ref objects, which reference specific strings within the protocol payloads. However, using the src/dst_payload_ref STIX objects for the TriStation protocols techniques, any detection utilizing these objects is likely to be fragile.
@@
Triconex (Table 3): Triton’s targeting of the Triconex safety controller required multiple adversary techniques used to manipulate devices that execute custom logic. Once the malicious program was downloaded to the device, it was executed on the device [T0821]. Then, the malware gained excessive privileges by executing a vulnerable system call on the device [T0890], and manipulated the running firmware of the device [T0857] to maintain an implant. Next, it enabled Command and Control (C2) [T0869] by manipulating a protocol handler to link to malicious code injected into the device [T0874]. Detecting this activity requires the ability to extract and audit programs on the device, such as through program upload functions or integrity checks. This could potentially be detected by the Process: Hash STIX object, while many devices produce CRC’s or checksums of deployed programs, these could be easily converted to a hash to support this object type. These Cyclic Redundancy Checks within the Triconex devices are a proprietary but documented function [triconex_safety_evaluation]. Most of these other Triconex behaviors are difficult to detect because they manipulate the device’s run-time memory, which depends on the undocumented and proprietary operation of the device. Detecting the exploitation of a vulnerable system call typically requires a runtime agent running on the device, while detecting the runtime manipulation in the underlying firmware requires mechanisms that perform integrity checks of a device’s OS and runtime environments. Therefore, these specific techniques likely cannot be detected through traditional detection platforms and STIX based information sharing.

### v2

@@ 1 Introduction
As critical infrastructure increasingly becomes the target of cyber threats, organizations and governments must be able to detect and share information about threat actor activity. Numerous government reports, such as Executive Order 13691 [14] and the Cyberspace Solarium Commission report [13], have highlighted the need for enhanced information sharing to safeguard critical infrastructure. Additionally, numerous government programs have been established to facilitate information sharing between the public and private sectors.
@@
In this paper, we examine the challenges that hinder the sharing of actionable threat information with asset owners within the ICS domain. This study leverages the MITRE ATT&CK for ICS [36] knowledge base, which aggregates cyber threat intelligence reports across known ICS incidents.
Our paper highlights the key challenges by reviewing three well-known incidents: Stuxnet [23], Industroyer [8], and Triton [29]. This analysis aims to identify the adversarial techniques employed by threat actors and to determine the information that asset owners need to implement effective detection mechanisms against these techniques. We also explore nine recent ICS vulnerability advisories with evidence of exploitation as documented in the CISA Known Known Exploitable Vulnerability (KEV) Catalog  [12].
@@ 2 Related Work
Recent advancements in Cyber Threat Intelligence (CTI) have underscored the complexities of securing ICS and broader Cyber-Physical Systems, particularly in the context of critical infrastructure. ICSs and CPSs often rely on proprietary and specialized protocols that are difficult to monitor and integrate with conventional cybersecurity tools. Krasznay et al. [18] highlighted the critical vulnerabilities in these systems, which stem from their original design priorities focused on safety rather than cybersecurity. The increasing connectivity of ICS and Operational Technology (OT) systems has exposed these vulnerabilities to sophisticated adversaries, including state-sponsored actors, necessitating a paradigm shift in cybersecurity practices. The SeConSys initiative in Hungary represents a significant milestone in addressing these challenges. This initiative facilitates sector-specific threat intelligence sharing and detection through the development of tailored CTI feeds, honeypots that simulate energy-sector environments, and the implementation of the Network and Information Systems (NIS) Directive. These efforts have been instrumental in promoting localized, collaborative approaches to cybersecurity and advancing frameworks for real-time threat information sharing among key stakeholders in the energy sector.
López-Morales et al. [19] extended these contributions by developing innovative methodologies that advance the state-of-the-art in CTI for CPSs. Their work introduces the $ICS^{2}$ Matrix, an enhanced version of the MITRE ATT&CK framework tailored specifically to ICSs. The $ICS^{2}$ Matrix incorporates novel attack techniques and mitigation strategies derived from an extensive analysis of Programmable Logic Controller (PLC)-related security research spanning 17 years. Additionally, the development of a high-interaction satellite honeypot provides a groundbreaking tool for simulating realistic adversary interactions, collecting data on emerging threats, and advancing the understanding of satellite-specific attack vectors. This work further proposes the creation of a sandbox environment for connected autonomous vehicles, enabling the simulation and analysis of cyberattacks on autonomous systems to improve resilience. Recently, MITRE extended its $D3FEND^{TM}$[22] cybersecurity ontology to OT, creating a structured knowledge base for defending cyber-physical systems, including new artifacts, such as controllers, sensors, and physical process actuators. Yet, their additional new artifacts and techniques failed to address structural issues in D3FEND’s core modeling approach, that were raised back in 2023[25], such as missing concepts, semantic overload, and a systematic lack of constraints that make the model under-specified. Although the new D3FEND extension represents significant practical improvements, OT users wan a primary representation for CTI exchange still need STIX, which D3FEND isn’t compatible with. Such contributions underscore the necessity of domain-specific CTI methodologies that address the unique operational and security challenges inherent in different CPSs.
@@ 2.1 Empirical Measurements and Quality Assessment of CTI Sharing
At the ecosystem level, empirical studies have begun to quantify the effectiveness and challenges of CTI sharing at scale. Jin et al. [17] provided one of the first comprehensive empirical measurements of CTI sharing effectiveness, analyzing STIX volume, timeliness, coverage, and quality across public sources. Their findings reveal significant variations in feed quality and highlight the challenges that organizations encounter while operating in a shared intelligence environment. This empirical foundation complements broader systematic reviews that examine how organizations actually use CTI for security decision-making [2] and identify persistent barriers to effective CTI sharing, including governance issues, misaligned incentives, and insufficient standardization [3].
Quality and trust of shared feeds remain persistent issues across the CTI ecosystem. Griffioen et al. [16] conducted systematic quality evaluations of cyber threat intelligence feeds, revealing inconsistent timeliness and specificity that can undermine the operational value of shared intelligence. Similarly, Schaberreiter et al. [31] developed quantitative frameworks for evaluating trust in CTI quality, demonstrating that trust and quality metrics vary significantly across different sources and feed types. Together, these studies highlight the need for more rigorous quality assurance mechanisms in CTI sharing initiatives.
@@ 2.2 Standardization and Technical Implementation
Standardization efforts play a crucial role in making threat observables portable and actionable across different organizational contexts. STIX 2.1, approved as an OASIS Standard (on 10 June 2021), serves as the exchange format used across many CTI producers and consumers [24]. However, challenges remain in effectively representing complex ICS-specific artifacts within existing STIX objects. Zych and Mavroeidis [39] have proposed enhancements to MITRE ATT&CK representations by STIX, to improve filtering and prioritization tasks, while RFC 9424 [26] provides important clarification on the capabilities and limitations of IOCs, reinforcing the value of higher-fidelity observables for creating durable detection capabilities.
@@
Recent advances in natural language processing have introduced new approaches for extracting threat intelligence from unstructured reports. While some pipelines pursue full STIX coverage through classical NLP techniques combined with external knowledge bases [20], alternative approaches rely on large language models to extract observables directly from text without requiring extensive knowledge graphs. These automated extraction techniques show promise for scaling CTI processing, though systematic evaluation reveals significant limitations and failure modes. Alam et al. [4] developed CTIBench, a comprehensive benchmark specifically designed to evaluate LLM performance in cyber threat intelligence tasks across multiple cognitive dimensions, including memorization, understanding, problem-solving, and reasoning. Their evaluation of five state-of-the-art models revealed concerning patterns: all models struggled significantly with tasks requiring domain expertise, with even the best-performing model (GPT-4) achieving only 71% accuracy on foundational CTI knowledge questions. More critically, their analysis identified systematic failure modes, including overestimation of threat severity scores, difficulty with out-of-knowledge cutoff information, and particularly poor performance on questions related to mitigation strategies and adversarial tools. These findings highlight substantial gaps in LLM reliability for cybersecurity applications.
@@
Fieblinger et al. [15] developed a comprehensive methodology for automating the extraction of actionable CTI using LLMs for generating Knowledge Graphs. Their approach systematically evaluates multiple open-source LLMs (including Llama 2 series, Mistral 7B Instruct, and Zephyr) across different extraction techniques, such as few-shot prompt engineering, guidance frameworks, and fine-tuning to optimize the extraction of meaningful triples from CTI texts. Notably, their methodology demonstrates that guidance frameworks significantly improve extraction performance beyond what is achievable with prompt engineering alone, while fine-tuning proves essential for extracting triples that adhere to specified ontologies. The extracted triples are subsequently used to construct knowledge graphs that provide structured, queryable representations of threat intelligence, facilitating downstream applications such as link prediction and threat analysis.
@@
The development of domain-specific evaluation datasets has become increasingly important for assessing automated threat intelligence processing. Bhusal et al. [7] developed the SECURE benchmark using MITRE ATT&CK as a foundation for generating cybersecurity LLM performance evaluation across extraction, understanding, and reasoning tasks. Their approach demonstrates the potential of leveraging structured threat intelligence frameworks like ATT&CK for creating systematic evaluation datasets. However, the reliability of LLM-based approaches in cybersecurity contexts remains a critical concern. Their comprehensive evaluation of seven state-of-the-art models revealed significant limitations, including hallucinations when models lack current information, difficulty with out-of-distribution detection, and substantial performance variations across different cybersecurity task types. Importantly, their findings demonstrate that while LLMs show promise for cybersecurity advisory roles, closed-source models consistently outperform open-source alternatives, particularly in problem-solving and out-of-distribution scenarios. These results underscore the importance of meticulous validation and human oversight when deploying LLMs for threat intelligence processing, particularly in critical infrastructure environments where inaccurate information can have severe consequences.
ATT&CK Framework Application in OT Environments.
@@
For technique-level grounding in OT environments, the MITRE ATT&CK framework and its ICS extension provide structured approaches to understanding adversary behaviors [35, 5]. Afenu et al. [1] have leveraged ATT&CK to validate ICS defense mechanisms and testbed implementations. Choi et al. [9] generated realistic OT attack sequences based on analysis of real-world datasets. These applications demonstrate the framework’s utility in bridging the gap between theoretical threat models and practical security implementations.
D3FEND Framework Extension for OT Environments.
TBD
@@
The existing literature highlights several recurring challenges in CTI, including limited interoperability of detection tools, insufficient standardization of threat-sharing protocols, and a lack of actionable intelligence for specific adversarial behaviors. While Krasznay et al. [18] emphasized the importance of legislative and collaborative frameworks, López-Morales et al. [19] focused on advancing technical methodologies to enhance the granularity and applicability of CTI. Both approaches converge on the need for tailored threat intelligence that integrates sector-specific requirements into practical and effective security solutions. Building on these foundational works, this study aims to address critical gaps in the ICS threat information-sharing ecosystem. Our research specifically focuses on expanding the cyber observable objects schema within the STIX standard to ensure the accurate representation of ICS-specific artifacts and adversarial techniques. Unlike prior studies that primarily concentrated on domain-specific or geographically localized advancements, this work seeks to generalize CTI enhancements across diverse ICS environments. By integrating detailed technical insights into standardized threat-sharing protocols, this study aspires to bridge the gap between theoretical threat models and actionable intelligence, thereby facilitating more effective collaboration and situational awareness across ICS stakeholders.
@@ 3 Key Roles and Functions in Sharing Threat/Vulnerability Information
Typically, governments facilitate information sharing by providing reports with threat or vulnerability information to critical infrastructures and other organizations [37]. The National Cyber Emergency Response Teams (CERT) is an important information-sharing hub, as it has unique access to threat information through the facilitation of information-sharing programs with both federal, defense, and intelligence agencies, as well as partnerships with private sector organizations.
Security vendors provide tools/platforms and information to support the detection and mitigation of cyber threats [34, 10]. Many of these vendors search for vulnerabilities and threat actor activity to develop analytics or rules to be incorporated into detection tools. Some vendors provide machine-readable threat information feeds, which can be directly consumed by security products, including those used to monitor for cyber threats. Further, many vendors obtain data from their detection infrastructures deployed at asset owner sites or collect information while supporting incident response efforts.
ICS vendors typically support information sharing efforts through the disclosure of security advisories and patches when a vulnerability is discovered in their software or devices [33]. Most importantly, ICS vendors have unique knowledge regarding the technical details for their products, which are typically based on proprietary technology.
Asset owners are responsible for managing the risk presented by the cyber threats. The larger organizations that maintain critical infrastructures typically have dedicated security operations teams, the smaller commercial asset owners commonly rely on their national CERT [27]. They may have a person partially employed in security operations to comply with regulations. The security team of an asset owner is responsible for monitoring their systems for threats and reporting to their auditors or the CERT.
@@
  • Standardized threat information languages: One key contribution to make threat information sharing more effective is the development of standard languages. For example, STIX standard provides a JSON format to communicate threat behaviors. This includes definitions of SCOs [24] for specific types of artifacts that can be used to detect malicious activities or collect information about them. Other formats include OpenIOC [^1] https://www.scribd.com/document/653654203/An-Introduction-to-OpenIOC and MISP XML [^2] https://www.misp-project.org/.
@@ 4.1.1 Triton
The Triton malware [11], also known as TRISIS or HatMan, targeted safety instrumented systems within an oil refinery by manipulating both an engineering workstation and the Triconex safety controllers used in the sulfur recovery units and burn management systems. While the broader campaign during this attack included activities used to gain access to the facility (references as TEMP.Veles in ATT&CK), the defined Tactics, Techniques and Proceduress specific to the ICS environment primarily address the behaviors associated with the manipulation of Windows-based engineering workstations, safety controllers, and the associated network communications between them. The following sections will review the adversary activity associated with these three areas based on the previously defined criteria.
@@
TriStation Protocol (Table 2): The TriStation protocol, which is a proprietary protocol without any available open-source parsers, was heavily used throughout the attack. The adversarial techniques using this protocol include discovering target devices [T0846], detecting and changing their operating mode [T0858] [T0868], downloading programs to the targeted safety controller [T0843], executing the program [T0871], and extracting programs from the device [T0845]. Developing detections for this activity requires the ability to (i) understand and parse the TriStation protocol  [21], (ii) identify specific functions of the protocol, and (iii) represent function payloads (e.g., program downloads contents). The STIX Network Traffic observable could be used to represent this activity, specifically the src/dst_payload_ref objects, which reference specific strings within the protocol payloads. However, using the src/dst_payload_ref STIX objects for the TriStation protocols techniques, any detection utilizing these objects is likely to be fragile.
@@
Triconex (Table 3): Triton’s targeting of the Triconex safety controller required multiple adversary techniques used to manipulate devices that execute custom logic. Once the malicious program was downloaded to the device, it was executed on the device [T0821]. Then, the malware gained excessive privileges by executing a vulnerable system call on the device [T0890], and manipulated the running firmware of the device [T0857] to maintain an implant. Next, it enabled Command and Control (C2) [T0869] by manipulating a protocol handler to link to malicious code injected into the device [T0874]. Detecting this activity requires the ability to extract and audit programs on the device, such as through program upload functions or integrity checks. This could potentially be detected by the Process: Hash STIX object, while many devices produce CRC’s or checksums of deployed programs, these could be easily converted to a hash to support this object type. These Cyclic Redundancy Checks within the Triconex devices are a proprietary but documented function [38]. Most of these other Triconex behaviors are difficult to detect because they manipulate the device’s run-time memory, which depends on the undocumented and proprietary operation of the device. Detecting the exploitation of a vulnerable system call typically requires a runtime agent running on the device, while detecting the runtime manipulation in the underlying firmware requires mechanisms that perform integrity checks of a device’s OS and runtime environments. Therefore, these specific techniques likely cannot be detected through traditional detection platforms and STIX based information sharing.